OpenStudy (anonymous):

If a > 0 and b < 0, then which of the following must be true. a + b < 0 a - b > 0 a + b > 0 ab > 0 b - a > 0

OpenStudy (anonymous):

a-b>0 since b is negative say b = -n, you have a-(-n)=a+n>0

OpenStudy (anonymous):

dang beat me to it xD